The radius of the sphere is ( 5.3 +- 0.1 ) cm the percentage error in its volume is ???

Grade:12th pass

1 Answers

Deepak Kumar Shringi
5 years ago
The volume is given as

V = (4/3)πr3

so,

V = (4/3) x 3.14 x 5.33

thus,

V = 623.30 cm3

now, the error equation

ΔV/V = 3(Δr/r)

or

ΔV = V x 3(Δr/r)

thus,

ΔV = 623.30 x 3x(0.1/5.3)

or error in surface area

ΔV = +/- 35.28 cm3

thus,

ΔV/V x 100 = (35.28/623.30) x 100

so,

ΔV/V x 100 = 5.66 %